Allison:

If YV= 28 in, Find the length of arc VYX

Allison:

1 attachment
Allison:

@dude

rishavraj:

Well you have the diamter, so you can the circumference right?? to get arc length you can use: \[(2~\pi~r)\times \frac{ \theta }{ 360 }\]

Allison:

Walk me through walk me throughhh

rishavraj:

https://youtu.be/tVcasOt55Lc check that out :)

jimthompson5910:

Minor arc UV is 117 degrees The central angle that cuts off this arc is also 117 degrees This is shown in red The adjacent blue angle adds with the red angle to get 180 (red angle) + (blue angle ) = 180 117 + B = 180 where B is the measure of the blue angle What is the value of B?

Allison:

78

jimthompson5910:

You'll need to subtract 117 from both sides of 117 + B = 180

jimthompson5910:

B = 78 is not correct

Allison:

I got it, I think it's 19.05

jimthompson5910:

That's not correct. It appears you may have been calculating the length of the orange arc (as shown below). I've marked this measure as 78 degrees This is because the blue arc is B = 180-117 = 63, and this 63 degree angle is congruent to its vertical angle counterpart (angle VZW) which adds to 15 to get 78 So angle VZX = 78 degrees

jimthompson5910:

But instead you want the arc as shown below. Note how it goes through points V, Y and X

jimthompson5910:

We can say (green arc) + (orange arc) = 360 green arc = 360 - (orange arc) green arc = 360 - 78 green arc = 282 degrees Use this to determine the length of the green arc

Allison:

My God, my brain

Allison:

Okay give me a while

jimthompson5910:

Its a lot to remember, yeah

jimthompson5910:

alright

Allison:

I'm so lost ):

jimthompson5910:

Do you see how to get the blue angle?

Allison:

No

jimthompson5910:

The red and blue angles must add to 180, since they form a straight line (red angle)+(blue angle) = 180 117 + B = 180 B = 180-117 B = 63 The blue angle is 63 degrees Agreed or no?

Allison:

Yes

jimthompson5910:

Do you also agree that the blue and orange angles are the same measure?

Allison:

Yes

jimthompson5910:

So because the orange angle is 63 degrees, it adds to the 15 degree angle to get 63+15 = 78 This is the measure of arc VWX Check out the diagram below

Allison:

Okay

jimthompson5910:

Since we want to deal with arc VYX, this means we want to find the length of the pink arc in the diagram below The pink and green arcs form a full circle. This means the two measures add to 360

Allison:

282

jimthompson5910:

correct, that's the measure of the pink arc in degrees

jimthompson5910:

This means (282/360) of the full circle is covered by the pink arc. Multiply that fraction by the circumference to get your answer

Allison:

68.90

jimthompson5910:

I'm getting (282/360)*pi*28 = 68.9055988687361 which rounds to 68.91

Allison:

Thank you <3

jimthompson5910:

You're welcome